OpenStudy (pounce0129):

how do you simplify 12n^5/36n

OpenStudy (zehanz):

Consider it two separate problems: \[\frac{ 12 }{ 36 }\]and\[\frac{ n^5 }{ n }\]I'm sure you know how to simplify the first one. For the second, remember n^5 = n*n*n*n*n. If you divide this by n, you loase one of these factors...